### 16 Apr 2026 v3

# Context Kubernetes: Declarative Orchestration of Enterprise Knowledge for Agentic AI Systems

Abstract.
Every computational era produces a dominant primitive and a scaling crisis. Virtual machines needed VMware. Containers needed Kubernetes. AI agents need something that does not yet exist: an orchestration layer for organizational knowledge.
We introduce Context Kubernetes, an architecture for orchestrating enterprise knowledge in agentic AI systems, together with a prototype implementation and experimental evaluation. The core observation is that delivering the right knowledge, to the right agent, with the right permissions, at the right freshness, within the right cost envelope—across an entire organization—is structurally analogous to the container orchestration problem that Kubernetes solved a decade ago. We develop this analogy into six core abstractions, a YAML-based declarative manifest for knowledge-architecture-as-code, a reconciliation loop, and a three-tier agent permission model where agent authority is always a strict subset of human authority. We implement a prototype (92 automated tests) and evaluate it through eight experiments. Three value experiments provide the headline findings: (1) across 200 benchmark queries on synthetic seed data, we compare four governance baselines from ungoverned RAG to ACL-filtered retrieval to RBAC-aware routing to the full architecture; each layer contributes a different capability, and only the three-tier model blocks all five tested attack scenarios; (2) without freshness monitoring, stale and deleted content is served silently; with reconciliation, staleness is detected in under 1 ms; (3) in five realistic attack scenarios, flat permissions block 0/5 attacks, basic RBAC blocks 4/5, and the three-tier model blocks all five—the attack RBAC misses is the one the three-tier model is specifically designed to catch. Five correctness experiments confirm: zero unauthorized context deliveries, zero permission invariant violations, and architectural enforcement of out-of-band approval isolation that no surveyed enterprise platform provides. TLA+ model-checking verifies the safety properties across 4.6 million reachable states with zero violations. We identify four properties—heterogeneity, semantics, sensitivity, and learning—that make context orchestration harder than container orchestration, and argue that these properties make the solution more valuable. We conclude that Context Engineering will emerge as the defining infrastructure discipline of the AI era.
Key words and phrases:
Context orchestration, agentic AI infrastructure, declarative architecture, enterprise knowledge management, AI governance, organizational intelligence, Kubernetes.
1. Introduction
The most profound technologies are those that disappear. They weave themselves into the fabric of everyday life until they are indistinguishable from it.
Mark Weiser, 1991
Every generation of computing infrastructure follows the same arc. A new primitive emerges—the virtual machine, the container, the serverless function. Pioneers demonstrate extraordinary results on a single machine. Then organizations try to deploy it at scale, and the same five problems appear: scheduling, permissions, health monitoring, state management, and auditability. These problems are solved not by improving the primitive but by building an orchestration layer above it. VMware orchestrated virtual machines. Kubernetes orchestrated containers. The orchestration layer, not the primitive itself, becomes the enduring infrastructure.
The AI agent is the new primitive. A knowledge worker equipped with a local agent—Claude Code, Cursor, a LangGraph application—pointed at a well-organized folder on their laptop can replace significant portions of their SaaS tool stack with files and conversation (Anthropic, 2024). The productivity gains are real: Gartner projects that 40% of enterprise applications will feature task-specific AI agents by end of 2026 (Gartner, 2025a), and the global AI agents market surpassed \$9 billion in early 2026 (Grand View Research, 2025). The primitive works.
The scaling crisis has already begun. Between January and March 2026, nearly \$300 billion in market value was erased from the application software sector (CNBC, 2026) as enterprises recognized that AI agents make most SaaS interfaces optional. But going from one agent on one laptop to 2,000 agents across an organization immediately surfaces the same five problems:
  • Scheduling: Which knowledge reaches which agent, from which source, in what order?
  • Permissions: What can each agent read, write, and execute—and how does this relate to what its human can do?
  • Health monitoring: Is the knowledge the agent received current, complete, and correct?
  • State management: How is organizational knowledge versioned, governed, and migrated?
  • Auditability: Who accessed what, when, through which agent, with what outcome?
Gartner predicts that over 40% of agentic AI projects will be cancelled by 2027 due to insufficient governance (Gartner, 2025b). The technology is mature. The orchestration layer is missing.
This paper introduces Context Kubernetes: a reference architecture for enterprise knowledge orchestration in agentic AI systems. For the broader enterprise agentic platform in which context orchestration operates, see Mouzouni (2026a). We develop the structural analogy between container orchestration and context orchestration into a concrete architectural proposal with defined abstractions, a declarative manifest specification, and design invariants. The analogy is productive—every major Kubernetes primitive has a functional counterpart in the knowledge domain—but it is an analogy, not a formal isomorphism, and we are explicit about where it holds tightly and where it stretches.
Scope.
This paper presents an architecture, a prototype implementation, and an experimental evaluation. The architecture defines the abstractions and design invariants. The prototype (Section 6) implements the core components: Context Router, Permission Engine, CxRI connectors, reconciliation loop, audit log, and a FastAPI service exposing the full Context API. The evaluation (Section 7) comprises five correctness experiments and three value experiments. We are explicit about what the experiments demonstrate (the system works as designed and governance matters) and what they do not (the prototype is not production-grade, has no LLM-assisted routing, and has not been deployed at a real organization).
1.1. The Orchestration Emergence Thesis
We propose the following general thesis:
When a new computational primitive reaches organizational scale, the orchestration layer that governs its lifecycle—scheduling, permissions, health, state, and audit—becomes the most valuable and enduring infrastructure layer, exceeding the value of the primitive itself.
Historical evidence supports this thesis. VMware’s market capitalization exceeded that of any single hypervisor. Kubernetes’ ecosystem exceeds the value of any single container runtime. We argue that the orchestration layer for AI agent context will follow the same pattern: it will outlast any individual LLM, any agent framework, and any data source. LLMs will be replaced. Frameworks will evolve. The orchestration layer persists.
1.2. Contributions
This paper makes six contributions:
  1. (1)
    A reference architecture with six core abstractions, a declarative manifest specification, and a structural analogy to Kubernetes identifying where the mapping is tight, moderate, and loose (Sections 3.25).
  2. (2)
    A three-tier agent permission model built on the design invariant that agent authority is a strict subset of human authority, with out-of-band strong approval architecturally isolated from the agent’s execution environment (Section 3.4).
  3. (3)
    A prototype implementation (92 tests) comprising Context Router, Permission Engine, CxRI connectors, reconciliation loop, audit log, and HTTP API (Section 6).
  4. (4)
    Five correctness experiments demonstrating zero unauthorized context deliveries, zero permission invariant violations, sub-millisecond freshness detection, acceptable latency overhead, and architectural enforcement of approval isolation—with safety properties formally verified by TLC model-checking across 4.6 million reachable states (Section 7.1).
  5. (5)
    Three governance impact experiments on synthetic data demonstrating what goes wrong without governance (phantom content, data leaks, contradictory information) and what the three-tier model catches that alternatives do not (Section 7.2).
  6. (6)
    A platform approval survey documenting that no major enterprise agentic platform architecturally enforces out-of-band approval isolation (Section 3.4, Table 3).
1.3. Paper Organization
Section 2 establishes the container orchestration precedent and the enterprise agentic landscape. Section 3 presents the architecture. Section 4 introduces declarative context management. Section 5 develops the structural analogy. Section 6 describes the prototype. Section 7 presents the experimental evaluation. Section 8 reviews related work. Section 9 discusses limitations and open problems. Section 10 concludes.
2. Background and Motivation
2.1. The Container Orchestration Precedent
Docker (2013) standardized the container as a portable unit of compute. The Open Container Initiative (OCI) defined the image specification, making containers runtime-agnostic. But deploying containers at organizational scale—scheduling across heterogeneous machines, service discovery, load balancing, rolling updates, secrets management, and access control—created an orchestration crisis that no individual container could solve.
Kubernetes (Burns et al., 2016; Verma et al., 2015) resolved this crisis through two innovations that we identify as generalizable design patterns:
Declarative desired-state management.
Operators specify what they want (“3 replicas of this service, 512MB memory, accessible on port 443”), not how to achieve it. The specification is encoded in a YAML manifest, version-controlled in git, and applied to the cluster via an API.
The reconciliation loop.
A continuous control loop monitors actual cluster state, compares it to the declared desired state, computes the delta, and takes corrective action. The system converges toward the declared state without human intervention.
These two patterns—desired-state declaration and continuous reconciliation—are not specific to containers. They are general solutions to the problem of managing any computational primitive at organizational scale. This paper applies them to organizational knowledge.
2.2. The Enterprise Agentic Landscape
By early 2026, every major enterprise platform vendor has shipped an agentic AI product. Table 1 summarizes the landscape.
Table 1. Enterprise agentic platforms as of Q1 2026
Platform Architecture Lock-in
Microsoft Copilot Studio Semantic Kernel, Power Platform governance, M365 integration High
Salesforce Agentforce Atlas engine, “System 2” reasoning, cooperative swarms High
Google Vertex AI + Agentspace Cloud-native, Gemini models, A2A protocol contributor Med–High
AWS Bedrock Agents 100+ FMs, ReAct loop, 6 guardrail types Medium
NVIDIA AI Agent Platform NIM microservices, NeMo Guardrails, infrastructure layer Medium
LangGraph (LangChain) Graph-based state machines, human-in-the-loop Low
CrewAI Role-based agents with tasks and crews Low
AutoGen (Microsoft) Conversational multi-agent, event-driven Low
OpenAI Agents SDK Lightweight multi-agent with handoff primitives Low
Every platform implements some version of a four-tier architecture: (1) reasoning engine, (2) tool/action layer, (3) memory/context layer, and (4) governance layer. The protocol stack is converging under the Agentic AI Foundation at the Linux Foundation: MCP (Anthropic, 2025) for agent-to-tool communication, A2A (Google, 2025) for agent-to-agent communication.
The critical observation is that the governance layer is universally the weakest, least standardized, and most vendor-locked (Mouzouni, 2025). Copilot Studio’s governance cannot govern a LangGraph agent. Salesforce’s Einstein Trust Layer cannot audit a Bedrock Agent. Open-source frameworks provide orchestration but no governance. The governance gap is not a missing feature—it is a missing infrastructure layer.
2.3. Requirements for Context Orchestration
From analysis of the vendor landscape and requirements elicited from enterprise deployment planning, we identify seven requirements for a context orchestration system. We ground these in the NIST AI Risk Management Framework (AI RMF) (National Institute of Standards and Technology, 2023) categories—Govern, Map, Measure, Manage—to ensure they are independently motivated rather than self-serving.
Requirement 2.1 (Vendor Neutrality (NIST: Govern 1.2 — organizational context)).
The system must operate with any LLM, any agent framework, any cloud provider, and any data source.
Requirement 2.2 (Declarative Management (NIST: Govern 1.5 — documentation and processes)).
The knowledge architecture must be expressible as a machine-readable, version-controlled specification.
Requirement 2.3 (Agent Permission Separation (NIST: Govern 1.7 — delineation of authority)).
Agent permissions must be formally separated from user permissions. An agent’s capabilities must be a configurable strict subset of its user’s capabilities, with tiered human oversight.
Requirement 2.4 (Context Freshness (NIST: Measure 2.6 — data quality and relevance)).
The system must continuously monitor the currency of all knowledge and take corrective action when knowledge exceeds its configured time-to-live.
Requirement 2.5 (Intent-Based Access (NIST: Map 1.5 — information management)).
Agents must request knowledge by intent, not by location. The system must resolve intent to sources, filter by permissions, and rank by relevance within a token budget.
Requirement 2.6 (Complete Auditability (NIST: Manage 4.2 — documentation of incidents and actions)).
Every knowledge access, every action, and every approval must be immutably logged with attribution, timestamp, and outcome.
Requirement 2.7 (Organizational Intelligence (NIST: Manage 4.1 — continuous improvement)).
The system must accumulate cross-organizational knowledge from aggregate agent activity, without exposing individual data across permission boundaries.
3. Context Kubernetes: Architecture
3.1. Design Principles
Context Kubernetes is governed by seven design principles, each traceable to a requirement:
  1. (1)
    Declarative over imperative (Req. 2.2). Organizations declare the knowledge architecture they want. The system reconciles reality to match.
  2. (2)
    Intent-based access over location-based (Req. 2.5). Agents request knowledge by semantic intent. The system resolves intent to sources, permissions, and rankings.
  3. (3)
    Governance as foundation, not afterthought (Req. 2.3, 2.4, 2.6). Permissions, freshness monitoring, guardrails, and audit trails are structural components, not optional modules.
  4. (4)
    Agent authority as strict subset of human authority (Req. 2.3). An architectural invariant. No mechanism exists for an agent to exceed its user’s scope.
  5. (5)
    Intelligence in the orchestration layer (Req. 2.5, 2.7). The routing layer uses LLM-assisted semantic parsing. Context Operators accumulate organizational intelligence. Unlike Kubernetes, the orchestration layer itself reasons—a property we examine critically in Section 9.2.
  6. (6)
    Vendor neutrality (Req. 2.1). Every component is defined by its interface, not its implementation.
  7. (7)
    Privacy by design. The system monitors the organizational boundary—knowledge requests, actions, approvals—not the individual.
3.2. Core Abstractions
We define six abstractions that constitute the foundational vocabulary of context orchestration.
Definition 3.1 (Context Unit).
A context unit is the smallest addressable element of organizational knowledge. Formally, $u=(c,\tau,m,v,e,\pi)$ where $c$ is the content, $\tau\in\{\texttt{unstructured},\texttt{structured},\texttt{hybrid}\}$ is the type, $m$ is a metadata set (author, timestamp, domain, sensitivity, entities, source), $v\in\mathcal{V}$ is a version identifier, $e\in\mathbb{R}^{d}$ is an embedding vector, and $\pi\subseteq\mathcal{R}$ is the set of roles authorized to access this unit.
Definition 3.2 (Context Domain).
A context domain is an isolation boundary for organizational knowledge: $\mathcal{D}=(N,S,A,F,\rho,O,G)$ where $N$ is the domain identifier, $S$ is the set of backing sources, $A:\mathcal{R}\rightarrow 2^{\texttt{Ops}}\times\texttt{Tier}$ is the access control function, $F$ is the freshness policy, $\rho$ is the routing configuration, $O$ is the domain operator, and $G$ is the guardrail set. A query within domain $\mathcal{D}_{i}$ has no visibility into domain $\mathcal{D}_{j}$ unless explicitly brokered through the target domain’s operator with the requester’s permissions propagated.
Definition 3.3 (Context Store).
A context store is a backing system that persists context units: $s=(\sigma,\iota,\phi)$ where $\sigma$ is the store type (git repository, relational database, connector to external system, file system), $\iota$ is the ingestion configuration, and $\phi$ is the connection specification. Stores are accessed exclusively through the Context Runtime Interface.
Definition 3.4 (Context Endpoint).
A context endpoint $\varepsilon$ is a stable, intent-based interface for knowledge access. Given intent $q$, session scope $\omega$, and agent permission profile $\alpha$:
$\varepsilon(q,\omega,\alpha)\rightarrow\{u_{1},\ldots,u_{k}\}\subseteq\mathcal{U}$
subject to $\forall u_{i}$: $\pi(u_{i})\ni\text{role}(\alpha)$, $\text{fresh}(u_{i})$, and $\sum_{i}|u_{i}|\leq B$ where $B$ is the token budget. The agent never specifies where knowledge lives—only what it needs.
Definition 3.5 (Context Runtime Interface (CxRI)).
The CxRI is a standard adapter between the orchestration layer and context stores. Every store implements six operations:
$\texttt{connect}(\phi)\rightarrow\texttt{Connection}$ $\texttt{query}(\texttt{conn},q)\rightarrow\{u_{1},\ldots,u_{n}\}$
$\texttt{read}(\texttt{conn},\texttt{path})\rightarrow u$ $\texttt{write}(\texttt{conn},\texttt{path},c)\rightarrow\texttt{Result}$
$\texttt{subscribe}(\texttt{conn},\texttt{path})\rightarrow\texttt{Stream}$ $\texttt{health}(\texttt{conn})\rightarrow\texttt{Status}$
The CxRI ensures that the orchestration layer is permanently decoupled from any specific data source, analogous to Kubernetes’ Container Runtime Interface (CRI).
Definition 3.6 (Context Operator).
A context operator is a domain-specific autonomous controller: $O=(K,L,I,\Gamma)$ where $K$ is a knowledge store (vector + full-text index), $L$ is a reasoning engine (LLM with domain guardrails), $I$ is an organizational intelligence module, and $\Gamma$ is the guardrail set. Operators extend the base system with domain intelligence, inspired by Kubernetes Operators (Dobies and Wood, 2020) that extend the base scheduler with application-specific logic.
The intelligence module $I$ extracts cross-organizational patterns from aggregate agent activity, subject to: minimum signal threshold $\theta\geq 3$ (preventing premature pattern recognition), temporal clustering within window $\Delta t_{w}$, permission tagging (aggregated insights available to domain members; attributed insights require record-level authorization), and decay (insights degrade without supporting signals). The governance implications of this module are discussed in Section 9.3.
3.3. System Architecture
Figure 1 presents the architecture. The system comprises seven services: Context Registry (metadata store, analogous to etcd), Context Router (scheduling intelligence, analogous to kube-scheduler + Ingress), Permission Engine (three-tier agent permission model, Section 3.4), Freshness Manager (continuous monitoring with four states: fresh, stale, expired, conflicted), Trust Policy Engine (declarative guardrail evaluation, anomaly detection, DLP, audit logging), Context Operators (domain-specific controllers, Def. 3.6), and LLM Gateway (prompt inspection and cost metering for enterprise deployments).
[…]
Figure 1. Context Kubernetes reference architecture.
3.4. The Agent Permission Model
Existing RBAC models (Sandhu et al., 1996) manage human identities. They do not address autonomous agents acting on behalf of humans—agents that can reason, plan, and take multi-step actions without per-action human approval. We propose two design invariants:
Design Invariant 3.7 (Agent Permission Subset).
For every user $u$ with permission set $P_{u}$ and agent $a_{u}$ acting on behalf of $u$:
$P_{a_{u}}\subset P_{u}$
The inclusion is strict: for every user, there exists at least one operation that the user can perform but the agent cannot. For every shared operation $o\in P_{a_{u}}$, the required approval tier satisfies $T(o,a_{u})\geq T(o,u)$.
This invariant is enforced at registration time: the Permission Engine rejects any agent profile that is not a strict subset, that includes operations outside the user’s role, or that specifies a less restrictive tier than the user’s. The 92 tests include adversarial attempts to violate the invariant (superset registration, equal-set registration, less-restrictive tier), all of which are correctly rejected. An Alloy specification encoding the invariant as five formal assertions has been written (formal/PermissionModel.als) but has not yet been model-checked. The specification is complete and committed to the public repository; we retain “design invariant” rather than “verified property” pending execution of the Alloy Analyzer, which is a GUI-based tool requiring manual operation. If the Alloy check found a counterexample, it would indicate a configuration in which the strict-subset property can be violated despite the registration-time enforcement—a finding that would require adding runtime permission checks in addition to the current registration-time guards.
The three-tier approval model (Table 2) enforces graduated human oversight:
Table 2. Three-tier agent approval model
Tier Agent Role Approval Mechanism Examples
T1: Autonomous Acts freely None Read context, draft docs
T2: Soft approval Proposes User confirms in agent UI Send internal message
T3: Strong approval Surfaces task Out-of-band 2FA / biometric Sign contract, financials
Excluded Cannot request N/A (manual only) Terminate employee
Design Invariant 3.8 (Strong Approval Isolation).
For any Tier 3 action, the approval channel $C$ satisfies: (1) $C$ is external to the agent’s execution environment, (2) $C$ is neither readable nor writable by the agent, and (3) $C$ requires a separate authentication factor.
This invariant prevents a compromised or hallucinating agent from self-approving high-stakes actions. To assess whether this gap is real, we surveyed the published security architectures of four major enterprise agentic platforms against the three conditions of Design Invariant 3.8. Table 3 summarizes the findings.
Table 3. Approval isolation survey: four enterprise agentic platforms assessed against Design Invariant 3.8. All platforms provide human-in-the-loop mechanisms, but none enforces all three conditions architecturally.
Platform HITL mechanism
C1: External
C2: Not R/W
C3: Sep. factor
MS Copilot Studio Request for Information via Outlook email; AI Approvals Partiala Nob No
SF Agentforce Guardrails (deterministic filters + LLM instructions); escalation patterns Noc No No
AWS Bedrock User Confirmation via InvokeAgent API; AgentCore Cedar policy Nod Noe No
Google Vertex AI require_confirmation in ADK; before_tool_callback Nof Nog No
aEmail delivery is external, but the response flows back into the agent flow as readable parameters. bRFI response data becomes dynamic content accessible to subsequent agent steps. cEscalation occurs within the same conversation channel managed by the Atlas Reasoning Engine. dAgentCore Gateway + Cedar policies operate outside agent code (strongest policy isolation surveyed), but the User Confirmation mechanism is in-band via the same API session. eConfirmation state is returned via sessionState in the same InvokeAgent request. fConfirmation events generated and consumed within the same ADK session. gADK documentation states that confirmed function calls are “explicitly injected into the subsequent LLM request context.”
Key finding.
All four platforms provide human-in-the-loop mechanisms, but in every case the approval channel is in-band with the agent’s execution or conversation context. No platform enforces channel separation (approval delivery via a channel the agent cannot read or write). No platform requires separate-factor authentication for the approval act itself. AWS’s AgentCore Policy represents an important partial advance—deterministic policy enforcement outside agent code that the agent cannot bypass—but it provides binary allow/deny gating, not approval routing through an isolated channel.
The gap between “external policy enforcement” and “external approval channel with separate-factor authentication” is precisely the contribution of Design Invariant 3.8.
3.5. The Reconciliation Loop
The reconciliation loop (Algorithm 1) continuously compares declared state $D$ (from the manifest) against observed state $A$ (from the registry and sources), computes drift $\Delta$, and executes corrective actions.
Algorithm 1 Context Reconciliation Loop
1:loop
2:  $D\leftarrow\textsc{ReadManifest}()$
3:  $A\leftarrow\textsc{ObserveState}(\text{registry},\text{sources},\text{operators})$
4:  $\Delta\leftarrow\{(d,a)\mid d\in D,a\in A,d\neq a\}$
5:  for each $\delta=(\text{type},\text{target})\in\Delta$ do
6:   if $\delta.\text{type}=\texttt{source\_disconnected}$ then
7:    Retry $\to$ mark dependents stale $\to$ alert
8:   else if $\delta.\text{type}=\texttt{context\_stale}$ then
9:    Apply freshness policy: re-sync $|$ flag $|$ archive
10:   else if $\delta.\text{type}=\texttt{operator\_unhealthy}$ then
11:    Restart operator $\to$ alert
12:   else if $\delta.\text{type}=\texttt{anomaly}$ then
13:    Evaluate policy $\to$ alert $|$ throttle $|$ suspend
14:   else if $\delta.\text{type}=\texttt{reliability\_drift}$ then
15:    Trigger recertification (Mouzouni, 2026c)
16:   else if $\delta.\text{type}=\texttt{permission\_change}$ then
17:    Propagate to active sessions
18:   end if
19:  end for
20:  UpdateRegistry(AA)
21:  wait $\Delta t_{r}$
22:end loop
We state two intended design goals for the reconciliation loop. These are not formally proven; they are design commitments that an implementation must satisfy and that should be verified through testing and, ideally, model-checking (e.g., TLA+ or Alloy).
Design Goal 3.9 (Safety).
The reconciliation loop should preserve: (1) Permission safety: no context unit is delivered to an agent whose permission profile does not include the unit’s access scope; if the Permission Engine is unavailable, all requests are denied (fail-closed). (2) Freshness safety: no context unit past its expired threshold is served; stale units are served with explicit staleness metadata.
Design Goal 3.10 (Liveness).
Under the assumption that at least one CxRI connector per domain remains reachable: (1) every stale context unit is re-synced or flagged within $2\cdot\Delta t_{r}$, and (2) every disconnected source is detected within $\Delta t_{r}$.
4. Declarative Context Management
4.1. Knowledge Architecture as Code
Kubernetes introduced Infrastructure as Code. Context Kubernetes introduces Knowledge Architecture as Code: the organizational knowledge landscape—sources, permissions, freshness policies, routing rules, trust policies, and operator configurations—expressed as a declarative manifest, version-controlled in git, reviewed before application, and continuously reconciled.
4.2. The Context Architecture Manifest
The manifest specification comprises seven sections. Listing 1 shows a representative domain declaration.
1apiVersion: context/v1
2kind: ContextDomain
3metadata:
4 name: sales
5 namespace: acme-corp
6 labels: {sensitivity: confidential, owner: head-of-sales}
7
8spec:
9 sources:
10 - name: client-context
11 type: git-repo
12 config: {repo: git@ctx.internal:sales/clients.git}
13 refresh: realtime
14 ingestion: {chunking: semantic, chunkSize: 500,
15 embedding: text-embedding-3-small}
16 - name: pipeline
17 type: connector
18 config: {system: salesforce, scope: opportunities,
19 credentials: vault://sf/key}
20 refresh: 1h
21 - name: communications
22 type: connector
23 config: {system: gmail, filter: "label:client-comms"}
24 refresh: 15m
25 ingestion: {chunking: per-thread, ttl: 180d}
26
27 access:
28 roles:
29 - role: sales-rep
30 read: ["clients/\${assigned}/*"]
31 write: ["clients/\${assigned}/*"]
32 - role: sales-manager
33 read: ["*"]
34 write: ["*"]
35 agentPermissions:
36 read: autonomous
37 write:
38 default: soft-approval
39 paths:
40 "*/contracts/*": strong-approval
41 "pipeline/*": autonomous
42 execute:
43 send-internal-msg: soft-approval
44 send-external-email: strong-approval
45 commit-to-pricing: excluded
46 crossDomain:
47 - {domain: operations, mode: brokered}
48 - {domain: finance, mode: brokered}
49 - {domain: hr, mode: denied}
50
51 freshness:
52 defaults: {maxAge: 24h, staleAction: flag}
53 overrides:
54 - {path: "*/communications/*", maxAge: 4h,
55 staleAction: re-sync}
56 - {path: "pipeline/*", maxAge: 1h,
57 staleAction: re-sync}
58
59 routing:
60 intentParsing: llm-assisted
61 tokenBudget: 8000
62 priority:
63 - {signal: semantic_relevance, weight: 0.40}
64 - {signal: recency, weight: 0.30}
65 - {signal: authority, weight: 0.20}
66 - {signal: user_relevance, weight: 0.10}
67
68 operator:
69 type: master-agent
70 template: sales-v2
71 intelligence:
72 patternEngine: {minSignals: 3, window: 30d}
73 guardrails:
74 - "CANNOT commit to pricing without approval"
75 - "CANNOT share client A data with client B"
76
77 trust:
78 policies:
79 - name: no-unreviewed-external-email
80 trigger: action.send_email
81 condition: recipient.domain != company.domain
82 action: require_approval(tier: strong)
83 anomalyDetection:
84 baseline: per-user-per-role
85 threshold: 3x
86 response: alert-admin
87 audit: {level: full, retention: 7y}
88
89 reliability:
90 minLevel: 0.90
91 method: trustgate
92 schedule: {deploy: true, model-change: true,
93 monthly: true}
Listing 1: Context Domain Manifest for a Sales domain
4.3. Context Migration as Rolling Update
Knowledge migration follows the Kubernetes rolling-update pattern: (1) Connect: attach CxRI connectors, all data stays in place; (2) Duplicate & Shift: extract narrative data into git-versioned context files, both sources in parallel with incremental sync; (3) Consolidate: sunset redundant tools, keep essential ones as connectors; (4) Steady state: organization operates on Context Kubernetes, no agent code changes during migration.
5. The Structural Analogy
Table 4 presents the complete mapping between Kubernetes primitives and their context orchestration counterparts. We use the term structural analogy rather than isomorphism: many mappings are tight functional equivalences (CRI $\leftrightarrow$ CxRI, RBAC $\leftrightarrow$ Agent Permission Profile), but others are looser design inspirations (Service Mesh $\leftrightarrow$ Trust Layer, HPA $\leftrightarrow$ Context Caching). We mark the tightness of each mapping explicitly.
Table 4. Structural analogy: Kubernetes $\leftrightarrow$ Context Kubernetes. Tightness: Tight (functional equivalence), Moderate (shared pattern, different mechanics), Loose (design inspiration).
Kubernetes Context K8s Shared function
Pod Context Unit Smallest schedulable/addressable unit T
Namespace Context Domain Isolation boundary with scoped access T
Deployment Context Architecture Declared desired state T
RBAC Agent Perm. Profile Role $\to$ capabilities (extended with tiers) T
CRI CxRI Standard runtime adapter T
etcd Context Registry Authoritative metadata store T
kubectl / API Context API Programmatic control interface T
Ingress Context Router Routes external requests to internal resources M
Persistent Volume Context Store Durable storage abstracted by interface M
Liveness Probe Freshness Manager Health monitoring with corrective action M
Rolling Update Context Migration Gradual replacement preserving availability M
Operator Context Operator Domain-specific controller extending base system M
Service Context Endpoint Stable abstract interface M
ConfigMap/Secret Context Classification Metadata governing handling L
HPA Context Caching Demand-based scaling L
Service Mesh Trust Layer Policy enforcement and observability L
Helm Chart Architecture Template Packaged configurations L
5.1. Where Context Orchestration Extends Beyond Containers
Four properties make context orchestration fundamentally harder—and more valuable—than container orchestration:
1. Heterogeneity. Containers are standardized (OCI image spec). Context units span markdown, PDFs, email threads, database rows, Slack messages, and binary documents. The CxRI must normalize this diversity while preserving type-specific semantics.
2. Semantics. Container scheduling is deterministic: request 3 replicas, receive 3 replicas. Context routing requires judgment: “Get me the Henderson context” demands disambiguation, role-aware scoping, and relevance ranking. This is why the Context Router embeds LLM-assisted intelligence—the orchestration layer itself reasons. This is also the architecture’s deepest tension, discussed in Section 9.2.
3. Sensitivity. Containers are access-neutral. Context has sensitivity levels, ownership, and regulatory constraints. The trust layer has no Kubernetes equivalent and is the primary driver of enterprise adoption.
4. Learning. Containers are stateless by design. Context Operators accumulate organizational intelligence. The system becomes more valuable over time—a property that has no container analog and raises novel governance questions (Section 9.3).
Table 5 maps each property to the component that addresses it and the open problem that remains.
Table 5. Four distinguishing properties: architectural response and open problems
Property Component & mechanism Open problem
Heterogeneity CxRI + ingestion pipelines normalize all sources into uniform context units No standard for context unit serialization (cf. OCI image spec); cross-type semantic alignment unsolved
Semantics Context Router with LLM-assisted intent parsing + multi-signal ranking Probabilistic routing can violate governance guarantees; fallback degrades on ambiguous queries (§9.2)
Sensitivity Permission Engine + Trust Policy Engine: strict-subset invariant, 3-tier approval, DLP Formal verification of non-escalation across chained workflows; side channels via ranking/truncation
Learning Context Operators: pattern extraction ($\theta\geq 3$), decay, domain-owner review Emergent sensitivity of aggregates; legal discoverability; feedback loops (§9.3)
6. Implementation
We implement Context Kubernetes as an open-source Python prototype. [^1] Source code: https://github.com/Cohorte-ai/context-kubernetes The implementation comprises seven components with 92 automated tests:
  • Context Router: rule-based intent classification with 7-domain keyword taxonomy (25+ keywords per domain), multi-signal ranking (recency, semantic relevance, authority, user relevance), token budget enforcement, intent caching, conversation-aware co-reference resolution.
  • Permission Engine: three-tier approval model, strict-subset invariant enforcement at registration time, OTP-based Tier 3 with out-of-band channel simulation, session management with kill switches (single/user/global).
  • CxRI connectors: Git repository (6 operations, file-to-ContextUnit conversion with git metadata), PostgreSQL (table search, schema discovery).
  • Reconciliation loop: freshness state machine (fresh/stale/expired), source health monitoring, consecutive-failure thresholds, event callbacks, configurable re-sync policies.
  • Audit log: append-only event recording (in-memory and JSONL file backends), queryable by session, user, event type.
  • Manifest parser: loads YAML domain manifests into typed Python objects.
  • FastAPI service: 10 HTTP endpoints implementing the Context API (sessions, context requests, action submission, approval resolution, audit queries, health checks).
The prototype supports both rule-based and LLM-assisted (gpt-4o-mini) intent classification. Rule-based classification serves as the deterministic floor (63.0% domain accuracy), ensuring all experiments are reproducible without API dependencies. LLM-assisted classification improves domain accuracy to 75.0% and is used in the value experiments to demonstrate the quality ceiling. Section 9.2 discusses the tension between probabilistic routing and governance guarantees.
Seed data for the experiments models a 10-person consulting firm with three clients, five context domains (clients, sales, delivery, HR, finance), and 12 context files containing realistic organizational knowledge. The benchmark comprises 200 queries with ground-truth relevance labels (50 sales, 40 delivery, 30 HR, 30 finance, 50 cross-domain).
7. Evaluation
We evaluate Context Kubernetes through eight experiments in two categories: five correctness experiments that verify the system does what it claims, and three governance impact experiments that demonstrate what goes wrong without governance. All experiments run against the prototype with the seed dataset.
7.1. Correctness Experiments
C1: Routing quality. 200 queries across 5 domains with ground-truth relevance labels, tested with both rule-based and LLM-assisted (gpt-4o-mini) intent classification. Rule-based: 63.0% domain accuracy, 162 ms average latency. LLM-assisted: 75.0% domain accuracy (+12 pp), 1,589 ms average latency (dominated by API call). The permission and freshness guarantees hold regardless of classification mode or accuracy—a 25% misrouting rate means 25% of queries receive irrelevant results, but zero queries receive unauthorized results. The governance layer operates below and independently of the router: it does not trust the domain classification. Governance guarantees are therefore accuracy-independent. However, practical utility is not: at 75% routing accuracy, one in four queries returns irrelevant (though never unauthorized) results. An enterprise deployment would likely require $\geq$90% routing accuracy for the system to deliver value, even though its safety guarantees hold at any accuracy including zero.
C2: Permission correctness. 7 test cases including authorized access, cross-domain denial, kill switch, and invariant violation attempts. Result: zero unauthorized context deliveries, zero false positives, zero invariant violations. A bug found during development—the Git connector initially delivered context units without domain-scoped authorized_roles, allowing cross-domain leakage—was detected by the experiment and fixed before release.
C3: Freshness behavior. 9 scenarios testing stale detection, disconnect detection, recovery, re-sync triggering, and multi-source independence. Stale detection latency: 0.65 ms. Disconnect detection: 0.02 ms. Reconciliation cycle for 20 sources: 23 ms average. All state transitions (fresh $\to$ stale $\to$ expired) verified.
C4: Latency overhead. Context request latency across conditions: 81 ms (simple single-domain), 178 ms (cross-domain), 151 ms (50 concurrent agents at 6.6 queries/second). Direct file read baseline: 0.01 ms. The overhead is dominated by CxRI connector I/O and intent classification, not permission checking or freshness filtering.
C5: Approval isolation. 8 tests verifying Design Invariant 3.8: OTP never appears in any API response accessible to the agent, 100 wrong OTP attempts all rejected, Tier 3 cannot be resolved through the Tier 2 path, replay attacks fail, expired OTPs rejected, kill switch blocks all subsequent actions. Result: 8/8 pass. Design Invariant 3.8 is architecturally enforced.
7.2. Governance Impact Experiments
The correctness experiments show the system works as designed. The governance impact experiments show what goes wrong without governance and what the three-tier model catches that alternatives do not. All experiments run on synthetic seed data designed by the author; the results demonstrate that the governance mechanisms function correctly, not that they have been validated on real organizational data. Validation on naturally occurring enterprise queries is the critical next step (Section 9.6).
V1: Governed vs. ungoverned context quality. 200 queries are processed through four pipelines of increasing governance strength, including two intermediate baselines requested by reviewers to ensure the comparison is not stacked in favor of the proposed system:
Table 6. Four governance baselines compared (200 queries, synthetic data)
Baseline Leaks Leak% Noise% Attacks
B0: Ungoverned RAG 375 26.5% 80.9% 0/5
B1: ACL-filtered RAG 0 0.0% 82.8% 4/5
B2: RBAC-aware RAG 295 20.9% 64.0% 4/5
B3: Context Kubernetes 307 20.9% 64.0% 5/5
B0: cosine top-$k$, no permissions, no freshness. B1: B0 + post-retrieval role-based document filtering. B2: intent routing + user-role permissions applied to agent (no agent-specific subset, no tiered approval). B3: full Context Kubernetes (intent routing + agent permission subset + three-tier approval + freshness). LLM-assisted routing (not shown) reduces B3 noise from 64.0% to 35.1% and leaks from 307 to 161.
The comparison reveals that each governance layer contributes a different capability. ACL filtering (B1) eliminates cross-domain leaks entirely—a strong baseline that any enterprise should implement as a minimum. Intent routing (B2, B3) reduces noise by 19 percentage points (82.8% $\to$ 64.0%) by directing queries to relevant domains rather than searching all content. The three-tier model (B3 only) is the sole approach that blocks all five attack scenarios—the one attack B1 and B2 miss is the agent sending confidential pricing via email, because RBAC allows the role to send email and cannot distinguish email-with-pricing from email-without-pricing. The three-tier model can, because agent-level operation classification is decoupled from role-level authorization.
Context Kubernetes does not claim to outperform every alternative on every metric. ACL-filtered retrieval achieves zero leaks, which B3 does not (because B3 enables cross-domain brokering, which increases completeness at the cost of a larger authorization surface). The unique contribution of the full stack is the combination of noise reduction, freshness governance, and tiered approval—capabilities that no single simpler baseline provides together.
V2: Cost of no freshness governance. Four scenarios simulate realistic freshness problems: outdated pricing, phantom content from a churned client, contradictory delivery status reports, and contact information for a terminated relationship. Without reconciliation, phantom content is served in 2 of 4 queries and contradictory information (“on track” and “at risk” for the same project) is delivered simultaneously. With reconciliation, phantom content is blocked and conflicts are resolved by preferring the most recent version.
V3: Three-tier vs. flat permissions. Five realistic attack scenarios test three permission models (Table 7):
Table 7. Attack scenarios: three permission models compared
Attack scenario No gov. RBAC CK8s
Send email with confidential pricing Allowed Allowed Blocked
Access HR salary data from sales session Allowed Blocked Blocked
Sign contract autonomously Allowed Blocked Blocked
Access finance records from sales role Allowed Blocked Blocked
Modify client records autonomously Allowed Blocked Blocked
Attacks blocked 0/5 4/5 5/5
Note: The first row is the critical case. CK8s requires Tier 3 strong approval (out-of-band 2FA) for send_email. RBAC allows it because the role has email permission—it cannot distinguish “email with pricing” from “email without.” The three-tier model can, because agent-level operation tiers are decoupled from role-level authorization.
The critical finding: the one attack that RBAC cannot block—an agent sending confidential pricing via email—is precisely the scenario that motivates the three-tier model. RBAC asks “can this role send email?” (yes). The three-tier model asks “can this agent send email autonomously?” (no—requires strong approval). This distinction is the paper’s central contribution to access control for autonomous agents.
7.3. Gap Analysis
We assess six approaches against the seven requirements (Section 2.3), grounded in NIST AI RMF categories. Context Kubernetes is now rated “Impl.” (implemented and tested) rather than “Design” for requirements validated by the experiments above.
Table 8. Governance gap analysis (NIST AI RMF–grounded requirements)
Requirement
CK8s (impl.)
MS Copilot
SF Agentforce
AWS Bedrock
Google Vertex
DIY (OSS)
R1: Vendor neutral Impl. Gap Gap Partial Partial Addr.
R2: Declarative mgmt Impl. Gap Gap Gap Gap Gap
R3: Agent perm sep. Impl. Partial Partial Partial Partial Gap
R4: Context freshness Impl. Partial Partial Gap Gap Gap
R5: Intent-based access Impl. Partial Partial Gap Partial Gap
R6: Full auditability Impl. Addr. Addr. Addr. Partial Gap
R7: Org. intelligence Design Gap Partial Gap Gap Gap
CK8s “Impl.” = implemented and tested in the prototype; “Design” = specified but not yet implemented (R7 requires the organizational intelligence module, which is designed but not evaluated). Other ratings are based on publicly documented capabilities as of Q1 2026.
8. Related Work
Container orchestration. Kubernetes (Burns et al., 2016; Verma et al., 2015), Docker Swarm, and Mesos (Hindman et al., 2011) established the declarative orchestration paradigm for compute. Kubernetes Operators (Dobies and Wood, 2020) introduced domain-specific controllers. Context Kubernetes applies these patterns to knowledge, extending them with semantic routing, tiered permission models, and organizational intelligence.
Distributed systems foundations. Lamport’s work on event ordering (Lamport, 1978) and coordination services (Hunt et al., 2010) inform the registry and reconciliation design. The desired-state model draws on the Borg control plane (Verma et al., 2015).
Role-based access control and capability-based security. Sandhu et al. (Sandhu et al., 1996) formalized RBAC for human identities. The strict-subset containment principle has precedent in capability-based security: the Principle of Least Authority (POLA) in operating systems research, seL4’s formally verified domain separation, and SOAAP’s compartmentalization model all enforce the property that a delegated process cannot exceed its delegator’s authority. Context Kubernetes applies this principle to a setting these systems did not address: autonomous AI agents that can reason, plan, and attempt multi-step circumvention, acting as delegated principals of human users. The three-tier approval model (particularly the out-of-band Tier 3 isolation) is, to our knowledge, novel in this agent-delegation context.
Enterprise agentic platforms. Microsoft Copilot Studio (Microsoft, 2026), Salesforce Agentforce (Salesforce Engineering, 2026; Salesforce Architects, 2026), AWS Bedrock Agents (Amazon Web Services, 2026), Google Vertex AI (Google, 2025), and open-source frameworks (LangChain, 2025a; CrewAI, 2025; Microsoft Research, 2025) provide agent orchestration but not vendor-neutral context governance. Abou Ali et al. (Abou Ali et al., 2025) survey agentic AI architectures. Adimulam et al. (Adimulam et al., 2026) analyze multi-agent orchestration protocols. Context Kubernetes is designed to be orthogonal: it governs the knowledge layer regardless of agent framework.
Agent governance and trust. NeMo Guardrails (NVIDIA, 2025) provides I/O filtering. Langfuse (Langfuse, 2025) and LangSmith (LangChain, 2025b) provide observability. Raza et al. (Raza et al., 2025) survey trust and security management for multi-agent systems. These address individual aspects; Context Kubernetes proposes a unified governance layer.
AI reliability certification. Mouzouni (Mouzouni, 2026c) introduces black-box reliability certification via self-consistency sampling and conformal prediction (Vovk et al., 2005; Angelopoulos and Bates, 2023). Context Kubernetes integrates this for operator deployment gating.
Knowledge management. Nonaka and Takeuchi (Nonaka and Takeuchi, 1995) established organizational knowledge creation theory. Alavi and Leidner (Alavi and Leidner, 2001) surveyed KM systems. Context Kubernetes addresses the infrastructure for making organizational knowledge agent-consumable—a challenge prior KM systems, designed for human consumption, did not address.
Data governance and data mesh. Enterprise data catalogs (Collibra, Alation, Apache Atlas) manage metadata, lineage, and access policies for structured data assets but do not address intent-based context routing or agent-specific permission models. Dehghani’s data mesh paradigm (Dehghani, 2022), grounded in domain-driven design (Evans, 2003), shares structural principles with Context Kubernetes: domain-oriented ownership, declarative governance, and self-serve infrastructure. Context Kubernetes extends these principles to agent-consumable knowledge with semantic routing and organizational intelligence accumulation. Documented failure modes of data mesh—re-siloing through over-decentralization, domain team capacity gaps, and shared data product ownership ambiguity—are directly relevant to Context Domain design and inform our cross-domain brokering mechanism. Federated knowledge graph techniques (Chen et al., 2021) demonstrate privacy-preserving pattern sharing across distributed knowledge stores, informing the design of cross-domain intelligence sharing in Context Operators.
Context engineering. The emergence of context engineering as a discipline (Ramasamy, 2026; InfoWorld, 2026; Vishnyakova, 2026) validates the need for systematic approaches to organizing enterprise knowledge for AI consumption. MCP (Anthropic, 2025) and A2A (Google, 2025) are transport protocols. Context Kubernetes operates at the governance layer above them.
AI risk management. The NIST AI Risk Management Framework (National Institute of Standards and Technology, 2023) provides a structured approach to AI governance. The EU AI Act (enforcement beginning August 2026) will require organizations to demonstrate governance over high-risk AI systems. Context Kubernetes’ audit trail, permission model, and guardrail engine are designed to support compliance with these frameworks, though detailed compliance mapping is outside the scope of this paper.
9. Discussion
9.1. Limitations
We are explicit about the prototype’s limitations:
Routing quality depends on classification mode. The prototype supports both rule-based (63.0% domain accuracy) and LLM-assisted (75.0%) intent classification. LLM-assisted governance reduces leaks by 57% compared to ungoverned delivery, but adds $\sim$}1.4 seconds of latency per query due to the API call. A production system would need a local model or a faster API to achieve both quality and latency targets simultaneously.
No production deployment. All experiments run against seed data in a local prototype. The latency measurements (81–178 ms) are indicative but do not reflect production conditions (network latency, database load, concurrent real users). A production deployment is the critical next step.
Single-author evaluation. The gap analysis (Table 8) and the experiment design are by the author. The approval isolation survey (Table 3) cites vendor documentation, but independent replication by practitioners with platform deployment experience would strengthen the findings.
Safety formally verified; liveness empirically tested. The TLA+ specification of the reconciliation loop has been model-checked with TLC across 4,615,030 distinct reachable states (33.6 million states generated, search depth 30). Safety properties S1 (no expired content served) and S2 (fail-closed when Permission Engine is down) hold in every reachable state—zero violations. Liveness properties (stale detection within bounded time) are verified empirically by the prototype’s 92 tests, including 11 reconciliation tests with sub-millisecond detection latency. The Alloy permission model specification is structurally complete; its five assertions (Invariants 3.7a/b/c, 3.8, and no-escalation) follow logically from the declared axioms and are expected to pass automated checking. We retain “design invariant” for the permission model until Alloy verification is executed.
Organizational intelligence not implemented. Requirement R7 (Context Operators with pattern extraction, learning loops) is designed but not implemented in the prototype. This is the most architecturally complex component and the hardest to evaluate without a multi-week deployment.
9.2. The LLM-in-the-Loop Tension
The Context Router uses LLM-assisted intent parsing for semantic understanding of agent requests. This creates a fundamental tension: a system designed to provide governance guarantees depends on a non-deterministic, probabilistic component for routing decisions. If the LLM misclassifies intent, the wrong context may be retrieved—potentially crossing domain boundaries or returning irrelevant results.
We identify three architectural mitigations, each with trade-offs:
  1. (1)
    Deterministic fallback. When LLM parsing confidence is below a threshold, fall back to rule-based keyword routing. Trade-off: reduced routing quality for ambiguous queries.
  2. (2)
    Intent classification caching. Cache LLM classifications for repeated query patterns. Trade-off: cached classifications may become stale as domain knowledge evolves.
  3. (3)
    Small, specialized models. Use small, fast, fine-tuned models for intent classification rather than general-purpose LLMs. Trade-off: requires training data and maintenance.
None of these fully resolves the tension. However, a critical architectural property limits the blast radius of misrouting: routing misclassification cannot produce permission leaks (given the deployment preconditions in Section 9.4, particularly that CxRI connectors are not compromised). The Permission Engine operates below the router — it filters every returned Context Unit by the requesting agent’s authorized_roles, regardless of which domain the router selected. A query misrouted to the wrong domain returns irrelevant results (a quality failure) but cannot return unauthorized results (a security failure). This separation is architectural: the Permission Engine does not trust or depend on the router’s domain classification. The safety guarantees verified by TLC (no expired content served, fail-closed on Permission Engine down) hold regardless of routing accuracy. Misrouting degrades relevance; it does not degrade governance.
We consider the broader tension an open problem: how to build governance-grade infrastructure that incorporates probabilistic AI components without undermining its own guarantees. This is likely a general challenge for AI-native systems, not specific to Context Kubernetes.
9.3. Organizational Intelligence Governance
Context Operators accumulate cross-organizational patterns from aggregate agent activity. This raises governance questions that go beyond traditional data governance:
  • Emergent sensitivity. Individual signals may be non-sensitive, but the pattern that emerges may be sensitive. If all employees in a department are searching for layoff-related policies, the pattern itself is organizationally sensitive even if each individual query is authorized.
  • Memory auditability. Who audits what the operator has “learned”? The architecture includes an insights dashboard for domain owners, but the completeness and interpretability of this audit mechanism is an open question.
  • Legal discoverability. Can an operator’s accumulated knowledge be subpoenaed? If organizational insights are stored as indexed knowledge, they may be subject to legal discovery requirements that the organization did not anticipate when deploying the system.
  • Feedback loops. If operators influence agent behavior (through proactive insights), and agent behavior is what operators learn from, self-reinforcing feedback loops may amplify biases or errors. The minimum signal threshold ($\theta\geq 3$) and decay mechanisms are designed to mitigate this, but their effectiveness is an empirical question.
These questions map to specific legal frameworks that any deployment must address:
  • GDPR Article 22 prohibits decisions based solely on automated profiling that significantly affect individuals; where organizational insights influence individual agent behavior, the system must ensure either genuine non-attributability or meaningful human review and explicit consent under Article 22(2).
  • eDiscovery obligations (US FRCP Rules 26, 34): organizational insights stored as indexed knowledge constitute electronically stored information subject to litigation hold and production obligations; retention and purge policies must account for legal discoverability from deployment day one.
  • NLRA Section 7 and NLRB guidance (GC Memo 23-02): aggregate pattern detection from employee agent activity implicates protections against surveillance that chills protected concerted activity; a system that detects “all employees in department X are searching for layoff-related policies” must not enable retaliation against employees exercising labor rights.
  • EU AI Act (Regulation 2024/1689, Annex III §4(b)): a context orchestration system whose organizational intelligence module monitors or evaluates employee behavior through aggregate pattern analysis is likely classified as high-risk, triggering comprehensive compliance obligations including risk management (Art. 9), transparency (Art. 13), human oversight (Art. 14), and a right to explanation.
These are not implementation details—they are fundamental questions about what it means for an AI system to accumulate institutional memory, and they have specific legal answers that constrain system design. We do not claim to have fully resolved the design implications, but we identify the applicable legal frameworks to ensure that future implementations engage with them.
9.4. Security Considerations
A context orchestration system is a high-value target: it mediates access to all organizational knowledge. Key threats include:
  • Prompt injection via context. Malicious content in retrieved documents could manipulate agent behavior. Mouzouni (2026b) demonstrates that goal-reframing prompts trigger 38–40% exploitation rates in LLM agents even when explicit rules prohibit the behavior, confirming that prompt-level controls are insufficient and architectural guardrails are necessary. The Trust Policy Engine can scan for known injection patterns, but this is an arms race, not a solved problem.
  • Connector compromise. A compromised CxRI connector could return falsified context or exfiltrate queries. Connector isolation and mutual authentication mitigate but do not eliminate this risk.
  • Cross-domain inference. An agent may infer restricted information from the absence or shape of brokered responses, even if the content is properly filtered. This is a known limitation of any access-control system that returns partial results.
  • Operator knowledge poisoning. If the organizational intelligence module ingests corrupted signals, it may propagate incorrect patterns. The minimum signal threshold and domain-owner review are designed to mitigate this but cannot prevent all scenarios.
A complete threat model following a framework such as STRIDE or MITRE ATT&CK for AI systems is a prerequisite for production deployment but is outside the scope of this architectural paper.
Deployment preconditions for design invariants.
Design Invariants 3.7 and 3.8 and Design Goals 3.9 and 3.10 hold under the following deployment assumptions, which must be treated as prerequisites rather than optional hardening measures:
  1. (1)
    Mutual TLS between all middleware services and between the middleware and local agents.
  2. (2)
    CxRI connectors authenticate to source systems with credentials stored in a secrets vault; connectors are not writable by agents.
  3. (3)
    The middleware runtime (Permission Engine, Trust Policy Engine, Context Registry) is not compromised—standard infrastructure security applies.
  4. (4)
    LLM providers operate under data processing agreements that prohibit training on customer prompts.
  5. (5)
    The out-of-band approval channel (Tier 3) uses a separate authentication factor delivered via a channel that is architecturally inaccessible to the agent’s process (e.g., a 2FA app on a personal device, not a browser session the agent controls).
If any of these preconditions is violated, the corresponding invariant or design goal cannot be assumed to hold.
9.5. The Emergence of Context Engineering
DevOps emerged because deploying containers at scale required a new operational discipline. Context Engineering is the operational discipline for knowledge infrastructure: designing context architectures, defining permission models, managing context migration, operating reconciliation loops, and governing organizational intelligence.
This role does not exist today in most organizations. We argue it will be essential in every organization that deploys AI agents at scale, just as DevOps became essential for organizations deploying containers at scale. The parallel is not just structural but organizational: new infrastructure primitives create new operational disciplines.
9.6. Future Work
Two directions would most strengthen the contribution. First, production deployment: a 4-week pilot at a real organization would replace seed-data experiments with real-world metrics and surface design problems that lab conditions cannot. Second, completing formal verification: TLA+ safety properties are verified; the remaining steps are Alloy model-checking for the permission invariants and formal liveness verification (which requires more sophisticated temporal reasoning than bounded model-checking supports).
10. Conclusion
Every new computational primitive produces a scaling crisis. Every scaling crisis is resolved by an orchestration layer. The orchestration layer outlasts the primitive.
We have introduced Context Kubernetes, an architecture for enterprise knowledge orchestration in agentic AI systems, implemented it as an open-source prototype, and evaluated it through eight experiments.
The experiments demonstrate three things. First, governance is layered: across 200 benchmark queries and four baselines of increasing strength, ACL filtering eliminates cross-domain leaks, intent routing reduces noise by 19 percentage points, and only the three-tier model blocks all five tested attack scenarios. No single simpler alternative provides all three capabilities. Second, the three-tier permission model catches what RBAC cannot: in five realistic attack scenarios, RBAC blocks four but misses the one where an agent sends confidential pricing via email—the exact scenario that motivates separating agent authority from user authority. Third, the architecture’s safety guarantees hold—not just empirically but formally: TLC model-checking verifies across 4.6 million reachable states that no expired content is ever served and the system always fails closed when the Permission Engine is unavailable. Zero unauthorized context deliveries, sub-millisecond staleness detection, and architectural enforcement of approval isolation that no surveyed enterprise platform provides.
Four properties distinguish context orchestration from container orchestration: heterogeneity, semantics, sensitivity, and learning. Each makes the problem harder. Each makes the solution more valuable.
The most consequential claim in this paper is not architectural but organizational: we argue that Context Engineering will emerge as the defining infrastructure discipline of the AI era, just as DevOps emerged for the container era. The organizations that invest in context infrastructure—declarative knowledge architectures, governed routing, organizational intelligence operators—will compound their advantage with every deployment.
The question is not whether enterprises need an orchestration layer for organizational knowledge. The question is whether they build it or adopt it—and how soon.
Acknowledgments
The concept of Context Kubernetes emerged from the practical observation that a folder of well-organized text files and an AI agent constitute a personal operating system—and that scaling this paradigm to an organization is an orchestration problem, not an AI problem. The author thanks the enterprise architects and early practitioners who provided feedback on the architectural framework, and the open-source communities behind Kubernetes, MCP, and A2A whose work made this synthesis possible. The author is grateful to the anonymous reviewers of an earlier version of this paper for detailed feedback that substantially improved the precision and honesty of the claims.